### v2

# A Physics-Inspired Optimizer: Velocity Regularized Adam

Abstract
We introduce Velocity-Regularized Adam (VRAdam), a physics-inspired optimizer for training deep neural networks that draws on ideas from quartic terms for kinetic energy with its stabilizing effects on various system dynamics. Previous algorithms, including the ubiquitous Adam, operate at the so-called adaptive edge of stability regime during training, leading to rapid oscillations and slowed convergence of loss. However, VRAdam adds a higher order penalty on the learning rate based on the velocity such that the algorithm automatically slows down whenever weight updates become large. In practice, we observe that the effective dynamic learning rate shrinks in high-velocity regimes, and damping oscillations. By combining this velocity‑based regularizer for global damping with Adam’s per‑parameter scaling, we create a powerful hybrid optimizer. For this optimizer, we provide rigorous theoretical analysis of operation at the edge of stability from a physical and control perspective for the momentum. Furthermore, we derive convergence bounds with the rate $\mathcal{O}(\ln(N)/\sqrt{N})$ for a stochastic non‑convex objective under mild assumptions. We demonstrate that VRAdam exceeds the performance against standard optimizers including AdamW. We benchmark various tasks such as image classification, language modeling, and generative modeling using diverse architectures and training methodologies including Convolutional Neural Networks (CNNs), Transformers, and GFlowNets.
1 Introduction
Optimizing the parameters of deep neural networks remains a cornerstone of progress in machine learning. Improving on the core idea of Stochastic Gradient Descent (SGD) Sutskever et al. (2013), adaptive methods like Adam (Adaptive Moment Estimation) Kingma & Ba (2017) have become ubiquitous due to their practical effectiveness across diverse tasks and architectures. Despite its success, the performance of Adam can be sensitive to hyperparameter choices and its training dynamics can exhibit instabilities Reddi et al. (2019). Furthermore, fully understanding the training dynamics of deep neural networks remains an open challenge Wang & Choromanska (2025), and even small improvements to existing optimization algorithms can often lead to significant reductions in resource consumption.
One line of work, observed empirically, is that training often occurs at the edge of stability Cohen et al. (2022; 2024), a regime for which the largest eigenvalue, also called sharpness, of the loss Hessian equilibrates around a fixed value proportional to the inverse of the learning rate (LR). This seems in contrast with common presumptions in classical optimization theory and has profound implications for convergence speed, stability, and generalization. In classical optimization, higher LRs lead to faster convergence at the cost of oscillations or divergence if stability constraints (depending on the loss landscape) are violated Boyd & Vandenberghe (2004). This pathological behavior of optimizers, like AdamW, leads to instabilities and a slowed decrease of the loss.
These challenges motivate the exploration of alternative optimization strategies. In line with the origins of machine learning itself Hopfield (1982); Ackley et al. (1985), one promising avenue draws inspiration from physics. For that, the optimization trajectory is conceptualized as a discretized motion of a particle within the high-dimensional loss landscape. Both from the structure of the “potential” landscape and the discretization, instabilities may arise from excessive “velocity” or overly large step sizes. This perspective suggests that mechanisms from high-energy and non-classical physics applied to optimization can improve aspects of stability. Building upon the established success of momentum, which incorporates velocity into gradient updates, recent ideas have explored a maximal velocity, drawing parallels to the speed of light in the theory of special relativity França et al. (2020).
[…]
Equipped with this new optimizer, VRAdam, we probe its dynamics at the adaptive edge of stability, observe faster convergence, and analyze its sharpness empirically against AdamW and Sharpness Aware Minimization (SAM) Foret et al. (2021). We also introduce rigorous theoretical analysis of the, global uniform exponential stability of momentum, Weber et al. (2024), as well as a physical inspired Lyapunov candidate derived from our Lagrangian that demonstrates stability properties. With well-tuned hyperparameters, we benchmark VRAdam against AdamW, RAdam Liu et al. (2020), SGD with Nesterov momentum and RMSProp Ruder (2017) on image classification with the CIFAR-10 dataset and the convolutional neural network architecture, on language modeling with transformers on the WikiText2 dataset, and a generative modeling task with GFlowNets and report improved performance on all tasks. We also report increased target-to-loss against AdamW on large scale training for language models such as GPT Brown et al. (2020) with marginal computational increase in overhead.
With this work, we contribute:
  • VRAdam, a physics-inspired and interpretable modification to AdamW,
  • Adaptive edge of stability analysis of VRAdam with faster convergence and associated empirical and theoretical evidence from momentum analysis,
  • Convergence bound for non-convex stochastic objective,
  • VRAdam outperforms AdamW and other optimizers on a wide range of benchmarks.
2 Background
Edge of stability. The training of deep neural networks does not follow classical optimization trajectories when trained with full-batch gradient descent (GD). However, during training, these models experience a surprising phase called the edge of stability (EoS). In this phase, the loss Hessian’s largest eigenvalue ($\lambda_{\text{max}}$) rises to approximately $2/\eta$, the numerical stability limit determined by the learning rate $\eta$. At EoS, the eigenvalue persists at this threshold, causing short-term, non-monotonic oscillations in the loss function. Despite these oscillations, the model still achieves long-term descent in the loss, though at the cost of slower convergence Arora et al. (2022); Cohen et al. (2022).
More recently, empirical bounds on the adaptive edge of stability (AEoS) have been observed for adaptive optimizers such as Adam Cohen et al. (2024). Here, the relevant stability threshold involves preconditioning the Hessian $H_{t}$, where the precondition is constructed from the exponential moving average (EMA) of past element-wise squared gradients $m_{t}$:
$P_{t}^{-1}H_{t},\quad P_{t}=\operatorname{diag}\left(\sqrt{m_{t}}+\varepsilon\right),$ (1)
This adaptive preconditioning coincides with the learning rate scaling in Adam (see Alg. 1) and scales down the step size in high‐variance (typically high‐curvature) directions as well as scales up in low‐variance ones. Since the local stability of an optimizer around a minimizer depends on the eigenvalues of the quadratic Taylor approximation $L(x)\approx\frac{1}{2}x^{\top}Hx$, Adam’s dynamics are shown to be stable Cohen et al. (2024) as long as
$\lambda_{\max}\left(P_{t}^{-1}H_{t}\right)<\frac{2+2\beta_{1}}{\left(1-\beta_{1}\right)\eta}=\frac{38}{\eta}\quad\left(\beta_{1}=0.9\right).$ (2)
However, as this threshold is attained, the adaptive oscillatory regime can slow final convergence, as the optimizer continually adjusts its preconditioner to maintain stability as well Song & Yun (2023).
Physical origins of exotic Lagrangians. To better comprehend and navigate the edge of stability regime, we can draw inspiration from the physics governing complex optimization scenarios. The deep insights provided by the interplay of physics and machine learning frameworks have been demonstrated in various scenarios, such as the improved interpretation of Neural Tangent Kernels (NTK) through Langevin dynamics Avidan et al. (2025). One central concept in physics is the Lagrangian $\mathcal{L}(x,v)=T(v)-V(x)$ of a system, which is a function of position $x$ and velocity $v$ and (typically) defined as the difference between kinetic energy $T$ and potential energy $V$ from which the equation of motion can be derived via the Euler-Lagrange equation. In this work, we investigate non-standard Lagrangian formulations of physical phenomena with excellent stability conditions. For example, the stability of seemingly disparate quantum systems like heavy quarkonia (described by NRQCD) and classical time crystals share conceptual parallels rooted in higher-order velocity terms. These terms fundamentally reshape energy landscapes by creating non-standard dispersion relations, establishing invariant submanifolds in phase space where stable configurations emerge as attractors or limit cycles. Further information regarding these systems are found in Appendix A.
3 Method
To translate this physics insight into an optimizer design, we identify the stabilizing aspects of such phenomena such as the heavy‐quark momentum with the optimizer’s global momentum buffer $v$. Accordingly, we posit a kinetic energy of the form: $T_{\mathrm{VRAdam}}(v)=\frac{m}{2}\|v\|^{2}+\frac{\beta_{3}}{4}\|v\|^{4}$, where $m$ is the mass and $\beta_{3}$ is a tunable parameter. The Lagrangian then becomes
$\mathcal{L}(x,v)=\frac{m}{2}v^{2}+\frac{\beta_{3}}{4}v^{4}-V(x),$ (3)
for which we solve the Euler-Lagrange equation $\frac{d}{dt}\frac{\partial\mathcal{L}}{\partial v}-\frac{\partial\mathcal{L}}{\partial x}=0$. We know that the loss of a neural network with parameters $x$ can be thought of as the potential landscape Holderrieth et al. (2024), such that $\frac{\partial\mathcal{L}}{\partial x}=-\frac{\partial V(x)}{\partial x}=-\nabla L_{\text{loss }}(x)$.
The Euler-Lagrange equation then becomes
$\frac{d}{dt}\left[\left(m+\beta_{3}\|v\|^{2}\right)v\right]=-\nabla L_{\text{Loss}}(x),$ (4)
which can be rearranged to
[…]
where the dot corresponds to the time derivative. Rather than explicitly constructing an optimizer based on an ordinary differential equation solver for Eq. 5 (via discretization and introduction of dissipation, e.g. França et al. (2020)), we utilize the term, $(1/(m+3\beta_{3}||v||^{2})$, and embed it as a dynamic learning rate into AdamW, to enhance the successfully proven properties. Note, that this expression from Eq. 5 is derived for the 1-D case, where the dynamics $\dot{v}$ is collinear with $v$. In particular, this approach avoids choosing specific forms associated with various different integrators. The full velocity regularized Adam is given in Alg. 1, where we highlight in blue the changes to AdamW.
In Fig. 1, the vector field in Eq. 5 is plotted for the case of the loss function being a simple quadratic. Compared to kinetic energy term without quartic velocity, the vector field is “squeezed” in $v$ direction, and the resulting trajectories are not circular. In this idealized setting, we can visualize the performance of VRAdam and Adam simplified to VRMomentum and Momentum, which corresponds to Alg. 1 without second-order moment estimates or bias corrections (setting $m_{t}=1$ and dropping lines 6, 8, and 9). After the first step, the lower step size of VRMomentum compared to Momentum can be seen, which leads to fewer oscillations.
Algorithm 1 VRAdam optimizer. $f(\theta)$: objective function; $\beta_{1},\beta_{2}\in[0,1)$; $v_{t}$: velocity estimate; $m_{t}$: second-moment estimate; $\eta_{t}$: dynamic learning rate at step $t$; $\alpha_{0}$: maximal learning rate; $\alpha_{0}/(1+\alpha_{1})$: minimal learning rate; $\beta_{3}$: velocity penalizer
1:Input: $f(\theta)$, $\theta_{0}$, $\alpha_{0}$, $\alpha_{1}$, $\beta_{1},\beta_{2}$, $\beta_{3}$, $\epsilon$, $\lambda$
2: Initialize $v_{0}\leftarrow 0$, $m_{0}\leftarrow 0$
3:for $t=1,\dots,T$ do
4:  $g_{t}\leftarrow\nabla f(\theta_{t-1})$
5:  $v_{t}\leftarrow\beta_{1}v_{t-1}+(1-\beta_{1})g_{t}$
6:  $m_{t}\leftarrow\beta_{2}m_{t-1}+(1-\beta_{2})g_{t}^{2}$
[…]
8:  $\hat{v}_{t}\leftarrow v_{t}/(1-\beta_{1}^{t})$
9:  $\hat{m}_{t}\leftarrow m_{t}/(1-\beta_{2}^{t})$
10:  $\theta_{t}\leftarrow\theta_{t-1}(1-{\color[rgb]{0,0,1}\eta_{t}}\lambda)-{\color[rgb]{0,0,1}\eta_{t}}\displaystyle\frac{\hat{v}_{t}}{\sqrt{\hat{m}_{t}}+\epsilon}$
11:end for
12:Output: $\theta_{T}$
Figure 1: Vector field $\dot{v}=-x/(1+3v^{2})$ and $\dot{x}=v$ derived by solving the Euler–Lagrange equation for $\mathcal{L}=v^{2}/2+v^{4}/4-x^{2}/2$. Black lines are continuous trajectories; blue/orange show VRMomentum vs. Momentum steps.
[…]
4 Analysis
4.1 Empirical Analysis
For this analysis, following Cohen et al. (2024), we train a ResNet 32 architecture on CIFAR-10 for an image classification task, with VRAdam, Adam and SAM. The training is stopped as soon as the training loss falls below 0.1 or an accuracy of 0.97 is reached. In Fig. 2 (a) and (b), the training curves of VRAdam indicate faster convergence in both minimizing training loss as well as maximizing training accuracy as compared to Adam and SAM, known for sharpness minimization. The training curves are smooth for both optimizers. When juxtaposed with the sharpness comparison depicted in Fig. 2 (c), we observe that the maximum eigenvalue (sharpness) of the preconditioned Hessian of the loss, remains adaptable to faster convergence due to the dynamic learning rate adjustments induced by VRAdam.
The effective learning rate of VRAdam is shown in Fig. 2 (d). During the first 25 iterations, the learning rate dynamically increases close to the maximal value allowed LR and then proceeds to decrease while exhibiting oscillatory behavior as we converge to the minima. The bound on the minimal LR is not active in this example, while the base LR of Adam stays constant throughout training. As described in the seminal work of Schmidhuber et al., we note that minima with lower sharpness are associated with better generalization Hochreiter & Schmidhuber (1997); Foret et al. (2021). We can also observe, that VRAdam’s dynamic learning rate quickly moves to the maximal LR to exploit the loss landscape optimally, while exploiting the trade-off between the adaptive edge of stability and faster convergence.
Figure 2: (a) Training loss curves for VRAdam, Adam, and SAM Foret et al. (2021) of ResNet 32 on CIFAR-10 (b) training accuracy curves (c) plot of maximal eigenvalues of the loss Hessian d effective learning rate during training. Hyperparameters for these plots are provided in Appendix D.4.
4.2 Stability of VRAdam
We analyze the behavior of VRAdam in the adaptive edge of stability regime compared to that of Adam in a momentum ablation setting without parameter-scaling based on second order moments estimates, bias corrections or weight decay. By simply adding weight decay, the admissible base step is constrained, the rest of the proof remains as follows.
Objective Function: Traditionally for stability analysis, we consider a convex quadratic objective:
$f(\theta)=\frac{1}{2}(\theta-\theta^{*})^{\top}H(\theta-\theta^{*})$
where $H$ is a positive definite Hessian matrix. Let $x_{t}=\theta_{t}-\theta^{*}$ be the error. The gradient is, $g_{t}=\nabla f(\theta_{t-1})=Hx_{t-1}$. Let $\mu=\lambda_{\min}(H)$ and $L:=\lambda_{\max}(H)$. From Algorithm 1, we know that $\eta_{t}\in\mathcal{H}:=\left[\eta_{\min},\alpha_{0}\right]$ and $\eta_{\min}=\alpha_{0}/\left(1+\alpha_{1}\right)>0$. Let $\beta_{1}$ from Algorithm 1 be $\beta$ for this analysis. The following theorem and proof provides sufficient but not necessary conditions for stability in discrete time. Thus proving a more restrictive case of the optimizer.
Theorem 4.1 (Uniform exponential stability of VRMomentum).
Consider $f(\theta)$ with $0\prec H\preceq LI$. Let VRMomentum be $\beta\in[0,1)$, $\beta_{3}>0$, $\alpha_{0}>0$, $\alpha_{1}\in(0,\infty]$, no bias correction, and $\lambda\geq 0$. If $\alpha_{0}L<B(\beta)=\frac{2(1+\beta)}{1-\beta}$ or $\eta_{\min}L<B(\beta)$ if the LR clip is active, then for any realization of $\{\eta_{t}\}$ generated by the gate $\eta_{t}=\alpha_{0}/(1+\min(\beta_{3}\|v_{t}\|^{2},\alpha_{1}))$, the origin is a globally uniformly exponentially stable equilibrium. Moreover, there exists a Common Quadratic Lyapunov Function (CQLF), $V(z)=z^{\top}Pz$ with $P\succ 0$ such that $V(z_{t})\leq(1-\epsilon)V(z_{t-1})$ for some $\epsilon\in(0,1)$.
Proof. Define the 2d-state $z_{t}=(x_{t},v_{t})$ and the parameterized update matrix
$A(\eta):=\begin{pmatrix}I-\eta(1-\beta)H&-\eta\beta I\\ (1-\beta)H&\beta I\end{pmatrix},\quad z_{t}=A(\eta_{t})z_{t-1}.$ (6)
Let $H=Q^{\top}\text{diag}(h_{i})Q$ with $0<\mu\leq h_{i}\leq L$. In the eigenbasis of $H$, the dynamics split into $d$ identical $2\times 2$ subsystems that share the same scalar $\eta_{t}$ and differ only by the curvature $h\in[\mu,L]$:
$\begin{pmatrix}\xi_{t}\\ v_{t}\end{pmatrix}=A_{h}(\eta_{t})\begin{pmatrix}\xi_{t-1}\\ v_{t-1}\end{pmatrix},\quad A_{h}(\eta):=\begin{pmatrix}1-\eta(1-\beta)h&-\eta\beta\\ (1-\beta)h&\beta\end{pmatrix}$. (7)
Hence it suffices to build a CQLF for the family $\{A_{h}(\eta):h\in[\mu,L],\eta\in\mathcal{H}\}$.
For a fixed $(h,\eta)$, the characteristic polynomial is
$\lambda^{2}-(1+\beta-\eta(1-\beta)h)\lambda+\beta=0,$ (8)
so the Schur criterion gives stability iff $\eta(1-\beta)h<2(1+\beta)$. In particular, if
$\alpha_{0}L<B:=\frac{2(1+\beta)}{1-\beta}$ (A1)
then every $A_{h}(\eta)$ with $\eta\in\mathcal{H}$ is Schur-stable. (This is the Adam/AEoS bound specialized to $P=I$ Cohen et al. (2024))
We now produce $P\succ 0$ such that
$A_{h}(\eta)^{\top}PA_{h}(\eta)-P\preceq-\epsilon I_{2}\quad\forall\eta\in\mathcal{H},\forall h\in[\mu,L],$ (A2)
for some $\epsilon>0$. Take the block-diagonal, curvature-agnostic form
$P=\text{diag}(p_{1}I,p_{2}I),\quad p_{1}>0,p_{2}>0,$ (A3)
which lifts to $P_{d}=\text{diag}(p_{1}I_{d},p_{2}I_{d})$ in $2d$ dimensions.
For a fixed $(h,\eta)$, set
$\Delta(\eta,h):=A_{h}(\eta)^{\top}PA_{h}(\eta)-P=\begin{pmatrix}\Delta_{11}&\Delta_{12}\\ \Delta_{12}&\Delta_{22}\end{pmatrix}$ (9)
with
$\Delta_{11}(\eta,h)$ $=p_{1}(-2\eta(1-\beta)h+\eta^{2}(1-\beta)^{2}h^{2})+p_{2}(1-\beta)^{2}h^{2},$ (10)
$\Delta_{22}(\eta,h)$ $=p_{1}\eta^{2}\beta^{2}+p_{2}(\beta^{2}-1),$ (11)
$\Delta_{12}(\eta,h)$ $=-p_{1}\eta\beta(1-\eta(1-\beta)h)+p_{2}\beta(1-\beta)h.$ (12)
We first force strict negativity on the diagonal, uniformly over $\eta\in\mathcal{H}$ and $h\in[\mu,L]$. Using the bounds $\eta\in[\eta_{\min},\alpha_{0}]$ and $h\in[\mu,L]$, a sufficient pair of conditions for $\Delta_{11}\leq-\delta_{1}$ and $\Delta_{22}\leq-\delta_{2}$ is
$\underbrace{(\alpha_{0}^{2}p_{1}+p_{2})L}_{\text{upper bound for the $h^{2}$-term}}$ $<\underbrace{\frac{2\eta_{\min}}{1-\beta}p_{1}}_{\text{lower bound for the linear $h$-term}},$ (A4)
$\underbrace{p_{2}(1-\beta^{2})}_{\text{negative term magnitude}}$ $>\underbrace{p_{1}\alpha_{0}^{2}\beta^{2}}_{\text{positive term bound}}.$ (A5)
(Here we used the worst cases $\eta=\alpha_{0}$ and $h=L$ wherever they make the expression largest.)
The pair (A4)–(A5) admits a solution $(p_{1},p_{2})$ iff
$\alpha_{0}^{2}L<\frac{2(1+\beta)}{\eta_{\min}}\quad\iff\quad\eta_{\min}>\frac{\alpha_{0}^{2}L}{2(1+\beta)}$ (A6)
Indeed, set $p_{1}=1$ and choose any
$p_{2}\in\left(\frac{\alpha_{0}^{2}\beta^{2}}{1-\beta^{2}},\frac{2\eta_{\min}}{(1-\beta)L}-\alpha_{0}^{2}\right).$ (A7)
The interval is nonempty precisely when (A6) holds.
Next, to pass from diagonal negativity to matrix negativity, we use the Schur complement:
$\Delta\preceq-\epsilon I_{2}\iff\Delta_{11}\leq-\epsilon,\quad\Delta_{22}-\frac{\Delta_{12}^{2}}{\Delta_{11}}\leq-\epsilon.$ (13)
Under (A4)–(A5), $\Delta_{11}$ and $\Delta_{22}$ are uniformly $\leq-\delta$ for some $\delta>0$. A direct (but routine) bound gives, for all $\eta\in\mathcal{H},h\in[\mu,L]$,
$\frac{\Delta_{12}^{2}}{-\Delta_{11}}\leq\frac{\beta^{2}(p_{1}\alpha_{0}+p_{2}(1-\beta)L)^{2}}{2\eta_{\min}p_{1}-(\alpha_{0}^{2}p_{1}+p_{2})(1-\beta)L},$ (14)
and the right-hand side is strictly smaller than $\Delta_{22}$ when (A4)–(A5) hold with slack. Thus there exists $\epsilon>0$ such that (A2) is satisfied.
For fixed $h$ and $P\succ 0$, the map
$\eta\mapsto A_{h}(\eta)^{\top}PA_{h}(\eta)=(A_{0}+\eta A_{1})^{\top}P(A_{0}+\eta A_{1})$ (15)
is matrix-convex in $\eta$ (its second derivative is $2A_{1}^{\top}PA_{1}\succeq 0$). Therefore, if the inequality
$A_{h}(\eta)^{\top}PA_{h}(\eta)-P\preceq-\epsilon I_{2}$ (16)
holds at the endpoints $\eta=\eta_{\min}$ and $\eta=\alpha_{0}$, it holds for all $\eta\in[\eta_{\min},\alpha_{0}]$. Under (A6), the choice (A7) does exactly that.
The block choice $P_{d}=\text{diag}(p_{1}I_{d},p_{2}I_{d})$ certifies
$A(\eta)^{\top}P_{d}A(\eta)-P_{d}\preceq-\epsilon I_{2d}\quad\forall\eta\in\mathcal{H},$ (17)
hence the quadratic Lyapunov $V(z)=z^{\top}P_{d}z$ yields
$V(z_{t})\leq(1-c)V(z_{t-1}),\quad c=\frac{\epsilon}{\lambda_{\max}(P_{d})}\in(0,1),$ (18)
which implies global uniform exponential stability and concludes the proof.
In the Appendix B, an alternative Lyapunov candidate based on Lagrangian physics is presented.
4.3 Convergence Analysis
In this section, we provide a convergence analysis of VRAdam by extending the work of Défossez et al. (2022) on Adam. The main result is a probabilistic bound that VRAdam converges in expectation to a stationary point for a nonconvex (and therefore also a convex) objective function under a set of regularity assumptions and simplifications.
Setup: The objective function is given by $F(\theta):\mathbb{R}^{d}\to\mathbb{R}$. $\nabla f$ are gradient estimates of $F$ such that $\mathbb{E}[\nabla f]=\nabla F$. The random variable $\tau_{N}$ describes an index with probabilities given by $\mathbb{P}[\tau=j]\propto 1-\beta_{1}^{N-j}$ for $j\in\{0,1,...,N-1\}$. The constants $\alpha_{\min}=\alpha_{0}/(1+\alpha_{1})>0$ and $\alpha_{\max}=\alpha_{0}>0$ correspond to the minimal and maximal possible learning rate at every step. $F$ is bounded from below, namely $\forall\theta,F(\theta)\geq F_{*}$. The gradient estimates are almost surely bounded in the $\ell_{\infty}$-norm, namely $\left\|\nabla f\right\|_{\infty}\leq R-\sqrt{\epsilon}$. $\nabla F$ is L-Lipschitz-continuous in the $\ell_{2}$-norm, namely $\forall\theta,\phi,\left\|\nabla F(\theta)-\nabla F(\phi)\right\|_{2}\leq L\left\|\theta-\phi\right\|_{2}$. The weight decay parameter $\lambda=0$ is set to zero. The bias correction for the velocity $\hat{v}_{t}\leftarrow v_{t}/(1-\beta_{1}^{t})$ is not included. The hyperparameter $\beta_{1},\beta_{2}$ fulfill $0<\beta_{2}<1$, $0\leq\beta_{1}<\beta_{2}$ and the number of iterates satisfies $N>\beta_{1}/(1-\beta_{1})$.
Theorem 4.2 (Convergence of VRAdam).
Given the six assumptions above, the iterates $\theta_{t}$ of Algorithm 1 fulfill:
$\mathbb{E}\bigl[\|\nabla F(\theta_{\tau})\|^{2}\bigr]\leq 2R\,\frac{F(\theta_{0})-F_{*}}{\alpha_{\min}\,\tilde{N}}\\ \quad+\,E\Bigl(\tfrac{1}{\tilde{N}}\ln\bigl(1+\tfrac{R^{2}}{(1-\beta_{2})\epsilon}\bigr)-\tfrac{N}{\tilde{N}}\ln(\beta_{2})\Bigr).$
with $\tilde{N}=N-\frac{\beta_{1}}{1-\beta_{1}},\quad\text{and}$
$E=\;\frac{\alpha_{\max}^{2}dRL(1-\beta_{1})}{\alpha_{\min}(1-\beta_{1}/\beta_{2})(1-\beta_{2})}+\;\frac{12\,\alpha_{\max}dR^{2}\sqrt{1-\beta_{1}}}{\alpha_{\min}(1-\beta_{1}/\beta_{2})^{3/2}\,\sqrt{1-\beta_{2}}}+\;\frac{2\,\alpha_{\max}^{3}dL^{2}\beta_{1}}{\alpha_{\min}(1-\beta_{1}/\beta_{2})\,(1-\beta_{2})^{3/2}}.$
A proof can be found in the appendix C. The right-hand side of the bound converges to zero in the limit of $N\to\infty$ if $\alpha_{min}(N)\propto N^{-1/2}$, $\alpha_{max}(N)\propto N^{-1/2}$, and $1-\beta_{2}(N)\propto N^{-1}$ to leading order. For this scaling of hyperparameters, VRAdam has the same convergence rate of $\mathcal{O}(\ln(N)/\sqrt{N})$ as Adam. While this convergence result provides a strong theoretical foundation, benchmarking is key to understanding performance under realistic conditions.
5 Benchmarks
We benchmark VRAdam against various optimizers on several datasets for various tasks with diverse architectures. These benchmarks include image classification using a CNN in the CIFAR-10 dataset Krizhevsky & Hinton (2009), language modeling with a Transformer architecture Vaswani et al. (2023) on the WikiText2 dataset Merity et al. (2016), and Generative Flow Networks (GFLowNets) Avidan et al. (2025), on a grid world for sequence generation tasks Chevalier-Boisvert et al. (2023), and a training a large language model (LLM) such as the Generative Pretrained Transformer (GPT) Brown et al. (2020). These benchmarks represent a broad variety of deep learning architectures and tasks, ranging from older techniques for image classification to much newer and actively emerging models and training techniques such as for GFLowNets. We focus on the comparison with AdamW since its exceeding popularity and consistent performance but also report the performance of stochastic gradient descent (SGD) with momentum Qian (1999), root mean square propagation (RMSProp) Ruder (2017) and RAdam Liu et al. (2020) for all tasks expect for the LLM benchmark. In particular, we demonstrate the excellent convergence properties as well as performance improvements in most benchmarks for VRAdam. Note that we run Bayesian optimization for hyperparameter sweeps for the image classification, and language modeling tasks, with several evaluations in close proximity to the found minimum, and the results are statistically significant. We report the optimal value from this optimization procedure, however, we include shaded error bars in the appendix. The Bayesian optimization for the hyperparameters was conducted with the objective to minimize validation loss, however, we also report test loss. Hyperparameters for the task involving GFlowNets were picked at random due compute constraints. All details regarding model training, setup and datasets along with error envelopes using multiple runs are provided in Appendices D and F.
LLM benchmark: In order to demonstrate the effectiveness of VRAdam for training large networks, such as large language models Brown et al. (2020), we present the validation loss results of VRAdam against that of AdamW when training a 124M parameter GPT-2 model on the FineWebEdu-10B dataset from scratch Penedo et al. (2024). We report the validation loss after around  57000 steps or 2 epochs as well as the average training time per epoch on a H100 GPU using hyperparameters found in the Appendix.
Method Training time (s) Validation Loss
AdamW 48549.56 3.511
VRAdam 48522.40 3.447
Table 1: Comparison of training time and validation loss for GPT-2 training.
Method WikiText-2 Loss CIFAR-10 Loss GridWorld Flow Matching Loss
Validation Test Validation Test Validation Test
VRAdam 5.99 6.00 0.476 0.469 1.25 1.33
AdamW 6.47 6.50 0.522 0.565 2.41 3.60
RAdam 7.511 7.554 2.300 4.005 1.407 2.290
SGD+Nesterov NaN NaN 0.625 0.620 2.71 2.61
RMSProp NaN NaN 0.801 0.813 25.0 25.0
Table 2: Comparison of optimizer performance across three tasks: language modeling on WikiText-2, image classification on CIFAR-10, and flow matching on GridWorld.
6 Related Work
Follow-up work on Adam The Adam optimizer was introduced in 2014 by Kingma & Ba (2017) by combining the previously existing concepts of momentum and scaling a base LR for each parameter based on second-order moment estimates. The base learning rate, however, remains hard-coded (potentially chosen through a learning rate scheduler) throughout training. Since then, several modifications to Adam have been introduced, such as NAdam Dozat (2016), RAdam Liu et al. (2020), Adabelief Zhuang et al. (2020), and in particular AdamW Loshchilov & Hutter (2017), which reintroduces weight decay in its original intention. LARS You et al. (2017) and LAMB You et al. (2020) compute learning rates for layers individually. More recent optimization techniques include LION, an automatically discovered alternative to signed momentum Chen et al. (2023), Sophia Liu et al. (2023), which uses estimated diagonal entries of the Hessian as a precondition, sharpness-aware minimization methods Foret et al. (2021), and a modified LR in Adam for reinforcement learning Ellis et al. (2024).
Understanding training dynamics, convergence analysis, and edge of stability Another line of work focuses on understanding the dynamics of the training of deep neural networks as well as derive convergence properties and guarantees for commonly used optimizers. Wang & Choromanska (2025) provides a recent survey over the later and Reddi et al. (2019) provide an explicit convex example for which Adam does not converge. A popularized framework for understanding training dynamics in the continuous training flow and infinite network width limit was introduced in Jacot et al. (2018), extended by finite width corrections in Huang & Yau (2020), and developed for graph neural networks in Du et al. (2019). Lastly, the role of the edge of stability regime offers an empirically view and was analyzed in Cohen et al. (2024); Arora et al. (2022); Cohen et al. (2022); Damian et al. (2022); Song & Yun (2023); Wang et al. (2022)
Symplectic Optimization This research area derives discrete‐time optimization algorithms by discretizing continuous‐time Hamiltonian or Lagrangian flows using symplectic integrators, which exactly preserve the underlying geometric (symplectic) structure of the dynamical system. This approach guarantees long‐term stability, energy‐preservation properties (or controlled energy dissipation in the dissipative case), and can provide valuable insights into existing optimizers. Notable work includes Betancourt et al. (2018); França et al. (2020); Maddison et al. (2018); Duruisseaux & Leok (2023); Yuan & Zhang (2023).
7 Discussion
Limitations We note that due to limited compute resources we were not able to benchmark the performance of VRAdam on model sizes beyond GPT 124M for a longer period. The edge of stability regime remains not fully understood and questions regarding generalization capabilities remain.
Summary of work Motivated by physical perspectives for complex optimization scenarios and stability conditions along with the adaptive edge of stability, we developed a new optimizer VRAdam based on quartic kinetic energy terms. We analyzed its performance at the adaptive edge of stability and benchmark it against several optimizers in particular AdamW on image classification, language modeling, and a generative task using GFlowNets where we report improved performance and robustness to hyperparameters.
Future work We hope that this work leads to further development of sophisticated and interpretable optimizers using concepts from physics, since even small improvement can lead to significant savings in resources.
Appendix A Bounding NRQCD
By explicitly breaking certain symmetries—Lorentz invariance in NRQCD and time-translation in time crystals— higher-order kinetic terms paradoxically enhance stability through topological protection mechanisms and the generation of emergent length/time scales Niemi (2021); Guha & Ghose-Choudhury (2019).
As a demonstration of this phenomenon, we can consider Nonrelativistic QCD (NRQCD), which is an effective field theory that expands full quantum chromodynamics in inverse powers of a heavy‐quark mass $m$ Assi et al. (2023). The bilinear heavy‐quark Lagrangian $\mathcal{L}$ up to $O(1/m^{3})$ reads,
$\mathcal{L}_{\mathrm{NRQCD}}=\psi^{\dagger}\left(iD_{0}-m+\frac{\mathbf{D}^{2}}{2m}+\frac{\mathbf{D}^{4}}{8m^{3}}+\cdots\right)\psi$ (19)
The corresponding Hamiltonian density is derived via Legendre transformation:
$\mathcal{H}_{\mathrm{NRQCD}}=\psi^{\dagger}\left(m-\frac{\mathbf{D}^{2}}{2m}-\frac{\mathbf{D}^{4}}{8m^{3}}-\cdots\right)\psi$ (20)
For a rigorous analysis of the dispersion relation, we work in momentum space:
$\psi(\vec{x},t)=\int\frac{d^{3}p}{(2\pi)^{3}}e^{i\vec{p}\cdot\vec{x}}\tilde{\psi}(\vec{p},t)$ (21)
In momentum space, the operators transform as:
$D_{0}$ $\rightarrow-iE\quad\text{(time evolution operator)}$ (22)
$\mathbf{D}$ $\rightarrow i\vec{p}\quad\text{(spatial covariant derivative)}$ (23)
$\mathbf{D}^{2}$ $\rightarrow-p^{2}\quad\text{(squared spatial derivative)}$ (24)
$\mathbf{D}^{4}$ $\rightarrow p^{4}\quad\text{(quartic spatial derivative)}$ (25)
The energy eigenvalue equation derived from the Hamiltonian is:
$E\tilde{\psi}(\vec{p})=\left(m+\frac{p^{2}}{2m}-\frac{p^{4}}{8m^{3}}+\mathcal{O}\left(\frac{1}{m^{5}}\right)\right)\tilde{\psi}(\vec{p})$ (26)
This gives the NRQCD dispersion relation to order $1/m^{3}$:
$E_{\text{NRQCD}}(p)=m+\frac{p^{2}}{2m}-\frac{p^{4}}{8m^{3}}+\mathcal{O}\left(\frac{1}{m^{5}}\right)$ (27)
A.1 Analysis of Boundedness
The exact relativistic energy-momentum relation:
$E_{\text{rel}}(p)=\sqrt{m^{2}+p^{2}}$ (28)
Using Taylor series expansion for $\sqrt{1+x}$ where $x=p^{2}/m^{2}$:
$\sqrt{1+x}=1+\frac{1}{2}x-\frac{1}{8}x^{2}+\frac{1}{16}x^{3}-\frac{5}{128}x^{4}+\mathcal{O}(x^{5})$ (29)
Applying this to the relativistic energy:
$E_{\text{rel}}(p)=m\left(1+\frac{1}{2}\frac{p^{2}}{m^{2}}-\frac{1}{8}\frac{p^{4}}{m^{4}}+\mathcal{O}\left(\frac{p^{6}}{m^{6}}\right)\right)$ (30)
Simplifying:
$E_{\text{rel}}(p)=m+\frac{p^{2}}{2m}-\frac{p^{4}}{8m^{3}}+\mathcal{O}\left(\frac{p^{6}}{m^{5}}\right)$ (31)
For the non-relativistic approximation without the quartic term:
$\lim_{p\to\infty}\frac{E_{\text{NR}}(p)}{E_{\text{rel}}(p)}$ $=\lim_{p\to\infty}\frac{m+\frac{p^{2}}{2m}}{\sqrt{m^{2}+p^{2}}}$ (32)
$=\lim_{p\to\infty}\frac{m+\frac{p^{2}}{2m}}{\frac{p}{\sqrt{1+\frac{m^{2}}{p^{2}}}}}$ (33)
$=\infty$ (34)
This shows that the non-relativistic approximation without the quartic term diverges from the true relativistic behavior.
The quartic term introduces a negative contribution to the energy that precisely cancels the fourth-order term in the relativistic expansion:
$E_{\text{NRQCD}}(p)=m+\frac{p^{2}}{2m}-\frac{p^{4}}{8m^{3}}+\mathcal{O}\left(\frac{1}{m^{5}}\right)$ (35)
Let’s define a parameter $\lambda=p^{2}/m^{2}$ (proportional to $v^{2}$). For the NRQCD expansion to be valid, we require $\lambda\ll 1$.
The second derivative of the NRQCD energy with respect to $\lambda$ is:
$\frac{d^{2}E_{\text{NRQCD}}}{d\lambda^{2}}=-\frac{3m}{4}\lambda+\mathcal{O}(\lambda^{2})$ (36)
This becomes negative for sufficiently large $\lambda$ within the domain of validity, indicating that the energy function develops a maximum rather than increasing unboundedly.
Appendix B Physics analysis
In this section, we provide a stability analysis from starting with a physical perspective, starting from our quartic Lagrangian from Eq. 3, we derive the 1D dynamics which we discretize implicitly through the state‑dependent step
$\eta_{t}=\frac{\alpha_{0}}{1+\min(\beta_{3}|v_{t}|^{2},\alpha_{1})},\qquad v_{t}=\beta_{1}v_{t-1}+(1-\beta_{1})\nabla V(x_{t-1}),\qquad x_{t}=x_{t-1}-\eta_{t}v_{t}.$ (37)
For a quadratic potential $V(x)=\tfrac{1}{2}x^{\top}Hx$ with $0\prec H\preceq LI$, we work in the eigenbasis $H=Q^{\top}\mathrm{diag}(h_{i})Q$. Writing coordinates as $(x_{i},v_{i})$ and $h\in(0,L]$, the one‑dimensional update reads
$v=\beta_{1}u+(1-\beta_{1})h,x,\qquad x^{+}=x-\eta v,\qquad y:=\eta h,\qquad\alpha:=1-\beta_{1},$ (38)
where $u:=v_{t-1}$ and $\eta\in[\eta_{\min},\alpha_{0}]$ with $\eta_{\min}=\alpha_{0}/(1+\alpha_{1})$. We introduce an energy that mirrors the quartic kinetic term and couples consistently to the curvature,
$\mathcal{E}h(x,v)=\frac{1}{2}h,x^{2}+\sigma,xv+\frac{\gamma}{2h},v^{2}+\frac{\delta}{4},v^{4},$ (39)
and set
$\sigma=\zeta,\frac{\beta_{1}}{2},\eta_{\min},\qquad\gamma=\zeta,\frac{\eta_{\min}}{1-\beta_{1}},\qquad\delta\geq\frac{\alpha_{0}^{2}}{\beta_{3}},\qquad\zeta\in(0,1).$ (40)
A direct expansion of $\mathcal{E}h(x^{+},v)-\mathcal{E}h(x,u)$ gives a quadratic form in $(x,u)$ plus the quartic difference $(v^{4}-u^{4})/4$. Because the coefficients of the quadratic form are convex polynomials of $y=\eta h$, the worst case over $y\in[\eta_{\min}h,\alpha_{0}h]$ occurs at the endpoints. Evaluating there, and using the adaptive edge‑of‑stability bound for momentum
$\alpha_{0}L<\frac{2(1+\beta_{1})}{1-\beta_{1}},$ (41)
one obtains constants $c_{x},c_{u}>0$ (depending only on $\beta_{1},\alpha_{0},\eta_{\min}$) such that, uniformly for all $h\in[\mu,L]$ and all $\eta\in[\eta_{\min},\alpha_{0}]$,
$\mathcal{E}h(x^{+},v)-\mathcal{E}h(x,u)\leq-,c_{x},h,x^{2}-,c_{u},\frac{u^{2}}{h}+\frac{\delta}{4},(v^{4}-u^{4}).$ (42)
The gate implies $\eta\leq\alpha_{0}/(1+\beta_{3}v^{2})$ and hence $\tfrac{1}{2}\eta^{2}v^{2}\leq\alpha_{0}^{2}/(8\beta_{3})$. The choice $\delta\geq\alpha_{0}^{2}/\beta_{3}$ therefore ensures that whenever $|v|\geq 1/\sqrt{\beta_{3}}$ the quartic contribution dominates any potential increase from the discretization term and contributes strict dissipation; when $|v|\leq 1/\sqrt{\beta_{3}}$ the negative quadratic terms already control the step. Summing over coordinates yields a global energy
$\mathcal{E}(x,v)=\sum{i=1}^{d}\mathcal{E}{h_{i}}(x_{i},v_{i}),$ (43)
and constants $\kappa_{1},\kappa_{2},\kappa_{4}>0$ such that
$\mathcal{E}(x_{t},v_{t})-\mathcal{E}(x_{t-1},v_{t-1})\leq-\kappa_{1}x_{t-1}^{\top}Hx_{t-1}-\kappa_{2}v_{t-1}^{\top}H^{-1}v_{t-1}-\kappa_{4}\sum_{i=1}^{d}\max\left\{0,v_{t,i}^{4}-\frac{1}{\beta_{3}^{2}}\right\}.$ (44)
This inequality formalizes the physical picture in Figure 1. The first two terms express curvature‑weighted exchange between potential and quadratic kinetic energy with net dissipation; the quartic term is a brake that activates precisely in high‑velocity regimes created near the adaptive edge of stability. The same mechanism explains the reduction in ringing and the lower sharpness observed in Figure 2: when $|v_{t}|$ grows, the gate reduces $\eta_{t}$ and the quartic channel increases dissipation, pushing the dynamics back to a low‑velocity regime. Two immediate consequences follow from the construction. First, the instantaneous adaptive stability threshold increases with the measured velocity:
$L_{\mathrm{EoS}}(t)=\frac{2(1+\beta_{1})}{(1-\beta_{1})\eta_{t}}=\frac{2(1+\beta_{1})}{(1-\beta_{1})\alpha_{0}}\left(1+\min(\beta_{3}|v_{t}|^{2},\alpha_{1})\right),$ (45)
so the method moves away from instability as oscillations grow. Second, each parameter update is uniformly bounded in norm by the gate,
$|x_{t}-x_{t-1}|=\eta_{t}|v_{t}|=\frac{\alpha_{0}|v_{t}|}{1+\beta_{3}|v_{t}|^{2}}\leq\frac{\alpha_{0}}{2\sqrt{\beta_{3}}},$ (46)
which prevents runaway steps and is not available to classical momentum. These properties are consistent with the design of Algorithm 1 and the empirical behavior reported in the analysis section.
Appendix C Convergence Proof for VRAdam
The proof is based on the work of the proof in Défossez et al. (2022).
There, the authors derive in equation A.37 the bound
$\underbrace{\tfrac{1}{2R}\sum_{n=1}^{N}\tfrac{\alpha_{n}}{\Omega_{n}}\sum_{k=0}^{n-1}\beta_{1}^{k}\,\mathbb{E}\!\left[\|G_{n-k}\|_{2}^{2}\right]}_{\mathbf{A}}$ $\;\leq\;F(x_{0})-F_{*}$
$\quad+\;\underbrace{\alpha_{N}^{2}L^{2}\sum_{n=1}^{N}\mathbb{E}\!\left[\|u_{n}\|_{2}^{2}\right]}_{\mathbf{B}}$
$\quad+\;\underbrace{\tfrac{\alpha_{N}^{3}L^{2}}{4R\sqrt{1-\beta_{1}}}\sum_{n=1}^{N}\sum_{l=1}^{n-1}\mathbb{E}\!\left[\|u_{n-l}\|_{2}^{2}\right]\sum_{k=l}^{n-1}\tfrac{\beta_{1}^{k}}{\sqrt{k}}}_{\mathbf{C}}$
$\quad+\;\underbrace{3\alpha_{N}R\sqrt{1-\beta_{1}}\sum_{n=1}^{N}\sum_{k=0}^{n-1}\!\Big(\tfrac{\beta_{1}}{\beta_{2}}\Big)^{k}\!\sqrt{k+1}\;\mathbb{E}\!\left[\|U_{n-k}\|_{2}^{2}\right]}_{\mathbf{D}}.$
By bounding $\alpha_{n}$ with $\alpha_{\min}$ in expression $A$, and $\alpha_{N}$ with $\alpha_{\max}$} in expressions $B,C,D$ we obtain the final result.
Appendix D Hyperparameters, Datasets and Model Architectures
Note, that we recommend $\beta_{3}=1.0$ as the default setting for VRAdam.
D.1 Image Classification
This section details the comprehensive sweep for the CNN, on the CIFAR-10 dataset.
  • Model: Convolutional Neural Network
  • Dataset: CIFAR-10
  • Hyperparameter sweep method: Bayesian optimization
  • Optimization metric: validation loss
Table 3: Dataset Splits during CNN Sweep
Dataset Split Configuration Description
Training 80% of the original CIFAR10 training set (50,000 images).
Validation 20% of the original CIFAR10 training set.
Test Full CIFAR10 test set (10,000 images).
Table 4: Fixed Hyperparameters during CNN Comprehensive Sweep
Parameter Value
Model Architecture Convolutional Neural Network
Dataset CIFAR-10
Epochs 100
Batch Size 1024
Scheduler Type (AdamW) WarmupCosineAnnealing
Warmup Epochs 5
Warmup Factor 0.1
Scheduler $\eta$ min $1\times 10^{-5}$
VRAdam $\beta_{1}$ 0.9
VRAdam $\beta_{2}$ 0.999
VRAdam power 2
VRAdam weight decay $1\times 10^{-5}$
VRAdam $\epsilon$ $1\times 10^{-8}$
AdamW $\beta_{1}$ 0.9
AdamW $\beta_{2}$ 0.999
AdamW weight decay $1\times 10^{-5}$
SGD momentum 0.9
SGD nesterov True
SGD weight decay $1\times 10^{-5}$
RMSProp $\alpha$ 0.99
RMSProp $1\times 10^{-8}$
RMSProp weight decay $1\times 10^{-5}$
Table 5: Swept Hyperparameters during CNN Comprehensive Sweep
Optimizer Parameter Sweep Configuration Optimal Parameter
VRAdam $\alpha_{0}$ Log‐uniform, Min: $1\times 10^{-4}$, Max: 0.1 0.0846
VRAdam $\beta_{3}$ Uniform, Min: 0.1, Max: 1.5 1.015
VRAdam $\alpha_{1}$ Integer Uniform, Min: 3, Max: 30 29
AdamW $\eta$ Log‐uniform, Min: $1\times 10^{-5}$, Max: $1\times 10^{-1}$ 0.0625
SGD $\eta$ Log‐uniform, Min: $1\times 10^{-5}$, Max: $1\times 10^{-1}$ 0.00784
RMSProp $\eta$ Log‐uniform, Min: $1\times 10^{-5}$, Max: 0.1 1.78e-4
D.2 Language Modeling
This section details the comprehensive sweep for the Transformer, on the WikiText-2 dataset.
  • Model: Transformer
  • Dataset: Wikitext-2
  • Hyperparameter sweep method: Bayesian optimization
  • Optimization metric: validation loss
Table 6: Dataset Splits during Transformer Comprehensive Sweep
Dataset Split Configuration Description
Training Full WikiText-2 predefined training set.
Validation Full WikiText-2 predefined validation set.
Test Full WikiText-2 predefined test set.
Table 7: Fixed Hyperparameters during Transformer Comprehensive Sweep
Parameter Value
Model Architecture TransformerModel
Epochs 100
Batch Size 32
Seed 0
Scheduler Type WarmupCosineAnnealing
Warmup Epochs 5
Warmup Factor 0.1
Scheduler $\eta$ $1\times 10^{-5}$
Model sequence length 64
Model embed dimension 128
Model hidden dimension 256
VRAdam $\beta_{1}$ 0.9
VRAdam $\beta_{2}$ 0.999
VRAdam power 2
VRAdam weight decay $1\times 10^{-5}$
VRAdam $\epsilon$ $1\times 10^{-8}$
AdamW $\beta_{1}$ 0.9
AdamW $\beta_{2}$ 0.999
AdamW weight decay $1\times 10^{-5}$
SGD sgd momentum 0.9
SGD sgd nesterov True
RMSProp $\alpha$ 0.1
RMSProp $\epsilon$ $1\times 10^{-8}$
Table 8: Swept Hyperparameters during Transformer Comprehensive Sweep
Optimizer Parameter Sweep Configuration Optimal Parameter
VRAdam $\alpha_{0}$ Log-uniform, Min: $1\times 10^{-5}$, Max: 0.1 1.55e-05
VRAdam $\beta_{3}$ Uniform, Min: 0.1, Max: 5.0 3.35
VRAdam normgrad Values: [True, False] False
VRAdam $\alpha_{1}$ Integer Uniform, Min: 5, Max: 30 7
Adam $\eta$ Log-uniform, Min: $1\times 10^{-5}$, Max: 0.1 1.661e-05
SGD $\eta$ Log-uniform, Min: $1\times 10^{-5}$, Max: $1\times 10^{-1}$ -
RMSProp $\eta$ Log-uniform, Min: $1\times 10^{-5}$, Max: 0.1 -
D.3 Generative Modeling with GFlowNets
Here we include the hyperparameters used for reporting the performance of the optimizers for the GFlowNet.
Table 9: Hyperparameters GFlowNets
Optimizer Parameter Value
VRAdam $\alpha_{0}$ 0.01
VRAdam $\beta_{3}$ 1
VRAdam normgrad False
VRAdam $\alpha_{1}$ 19
VRAdam weight decay $1\times 10^{-5}$
AdamW weight decay $1\times 10^{-5}$
AdamW $\eta$ 0.01
SGD $\eta$ 0.01
RMSProp $\eta$ 0.01
D.4 Edge of Stability analysis
Table 10: Hyperparameters edge of stability analysis
Parameter Value
Model Architecture ResNet 32
Max iterations 20000
Batch Size 1000
Seed 0
Loss criterion Mean squared error
VRAdam $\alpha_{0}$ 0.002
VRAdam $\beta_{1}$ 0.9
VRAdam $\beta_{2}$ 0.999
VRAdam $\beta_{3}$ 1
VRAdam power 2
VRAdam normgrad False
VRAdam $\alpha_{1}$ 19
VRAdam $\epsilon$ $1\times 10^{-7}$
Adam $\beta_{1}$ 0.9
Adam $\beta_{2}$ 0.999
Adam $\epsilon$ $1\times 10^{-7}$
Appendix E Compute Resources
All experiments were run on Lambda cloud instances or the Google cloud platform (GCP). Experiments were conducted either using a NVIDIA L4 GPU with 24 GB of GPU memory and 31 GB of system memory or larger experiments were performed on a NVIDIA A10 with 24 GB of GPU memory and 200 GB of system memory. The GPT benchmark was run on an NVIDIA H100 GPU.
Appendix F Loss curves with Error Envelopes
Figure 3: Train (left) and validation (right) loss curves with error envelopes calculated using different run values for language modeling using AdamW.
Figure 4: Train (left) and validation (right) loss curves with error envelopes calculated using different run values for language modeling using VRAdam.
Figure 5: Train (left) and validation (right) loss curves calculated using different run values for language modeling using SGD Nesterov with momentum. The dots on the top indicate NaN values.
Figure 6: Train (left) and validation (right) loss curves with error envelopes calculated using different run values for language modeling using RMSProp. The dots on the top indicate NaN values.
Figure 7: Train (right) curves with error envelopes calculated using different run values for image classification using VRAdam. Test loss values (left) over different runs.
Figure 8: Train (right) curves with error envelopes calculated using different run values for image classification using AdamW. Test loss values (left) over different runs.
Figure 9: Train (right) curves with error envelopes calculated using different run values for image classification using SGD Nesterov with Momentum. Test loss values (left) over different runs.
Figure 10: Train (right) curves with error envelopes calculated using different run values for image classification using RMSProp. Test loss values (left) over different runs.